OpenStudy (anonymous):

Michael is 51 years old and Vanessa is 9 years old. How many years will it take until Michael is only 4 times as old as Vanessa?

OpenStudy (anonymous):

\[M = 51 + n, V = 9 + n\]\[4V = M\] Now find the n. You know how to?

OpenStudy (anonymous):

thank you very much :)

OpenStudy (anonymous):

yes i do